Determine the price p which result in maximum total revenue
1
Expert's answer
2021-09-01T17:57:48-0400

"p=500-\frac{1}{50}q,\\ \text{where p=price and q=demand}\\
R(q)=p×q=500q-\frac{1}{50}q^2, text{where R = revenue}\\
\text{And, maximum revenue is given by derivative of Revenue function}\\
R'(q)=500-\frac{2}{50}q\\
0=500-\frac{2}{50}q\\
q=12500\\
p=250
\text{Therefore, Revenue R is maximum when price p=250.}"